# RAVEN: Realtime Accessibility in Virtual ENvironments for Blind and Low-Vision People

### v1

@@ 4.3.2. Procedure
Studies lasted approximately 1.5 hours and were conducted either in-lab (N=6) or at participants’ homes (N=2). Participants used a researcher’s device running Unity, a full-sized keyboard, and stereo headphones. Each session followed the three scene explorations (section 4.2), then a survey (usefulness of categories, confidence, intuitiveness, and SUS (Brooke, 1996)), and finally a semi-structured interview (usability, new category ideas, prompt strategies, accessibility needs, AI trust). Sessions were audio- and video-recorded. Participants received a \$50 gift card and travel reimbursement. The study was IRB-approved.
@@ 4.3.3. Data Analysis
We analyzed three sources of data: survey ratings, prompt logs, and interview transcripts. Together, these provided complementary perspectives on participants’ experiences and interaction strategies.
@@
We computed descriptive statistics (means, standard deviations) for ratings of category usefulness (section 4.1), system intuitiveness, user confidence, and SUS scores (Brooke, 1996) to assess perceived usability.
@@
Prompts and system responses from Scenes 2 and 3 were recorded, yielding 336 valid prompts (181 from Scene 2 and 155 from Scene 3), after excluding the tutorial scene and erroneous inputs. Each prompt was independently coded by multiple researchers along three dimensions, with disagreements resolved through discussion:
@@
Interviews were transcribed and analyzed using applied thematic analysis (Guest et al., 2012). Three researchers collaboratively: (1) familiarized themselves with the transcripts, (2) tagged data with codes, (3) developed an initial codebook, (4) refined codes and themes, (5) finalized the codebook, and (6) synthesized themes into findings. This process produced a final codebook with 114 codes, organized into 30 third-level themes, 12 second-level themes, and 4 first-level themes.
@@ 5. User Study Findings
Our analysis highlights both the promise and current limitations of RAVEN. We first report participants’ overall perceptions of usability, then examine performance across the six guiding and four emerging prompt categories, and finally describe the strategies and goals that shaped how participants engaged with the system.
@@ 5.1. Overall System Usability
Despite these advantages, several limitations emerged. Of the 336 valid prompts, 253 (75.3%) produced a correct query answer or intended modification (determined through researcher coding of session recordings), 9 (2.7%) were correctly flagged as out-of-scope, and 74 (22.0%) failed. Failures included intent errors (14 prompts, where the LLM misunderstood participant goals) and technical errors (62 prompts, where the LLM hallucinated, misrepresented system capabilities, or failed in code execution). P4-LP described these as gaps in flexible language and cross-modal consistency: extinguishing a torch stopped the light but left the fire sound. P6-VI similarly noted misapplied assumptions, such as turning all objects white after asking to “make the room white”. Participants (N=7) raised trust concerns about misleading responses. As P2-VI asked, “If it gives the wrong information, how is a visually-impaired person going to know?” P4-LP concluded: “right now, [the system] is not so foolproof as to say I completely trust it.” These findings highlight both the promise and current fragility of the approach.
5.2. Performance Across Prompt Categories
Qualitative coding of 336 prompts revealed ten categories: six guiding categories defined in study design (Object Location, Audio Volume, Color, Object Size, Scene Brightness, Audio Pitch) and four emerging categories that surfaced during analysis (Scene Description, Semantic Description, Functionality, Creation/Deletion). Nine prompts were coded as “Other” and eight combined multiple categories (e.g., , “Make the bench normal-sized (Object Size) and put us on the bench” (Object Location)). See table 3 for code definitions and examples.
@@ 5.2. Performance Across Prompt Categories
(a) Occurrences of each category (blue = guiding, orange = emerging).:
(b) Correctness breakdown per category.:
@@ 5.2.1. Guiding Categories
Lowest rated ($M=2.8,SD=1.7$), Audio Pitch adjustments were less useful overall. Some participants envisioned benefits for distinguishing status (e.g., damage to a game character) or orientation (P7-B: “using pitch for positional information is a really good idea… for example, gets higher when you get closer to it”), but most preferred volume as a more salient cue. This category revealed potential for system-level pitch cues for important events or orientation, but limited value as a user-driven adjustment.
@@ 5.2.2. Emerging Categories
Although not predefined, these categories were frequently used, revealing additional details preferred by participants. Scene descriptions offered general overviews, while semantic descriptions probed higher-level qualities (e.g., , the mood of a cat or the content of a laptop screen). Participants valued detail but noted issues of verbosity: P4-LP suggested limiting output to nearby objects or allowing control over granularity. P7-B emphasized context-aware tailoring: “Maybe all I need to know is… I don’t need to know that it’s a Toshiba.” Out-of-scope acknowledgments were common, as users often sought knowledge beyond system capabilities. Suggested improvements included adding shape (P1-LP), clearance (P7-B), sound characteristics (P2-VI, P8-B), and hazard indicators (P7-B).
@@
Although infrequent, these categories revealed unmet needs. Participants wanted to interact with objects functionally (e.g., , “sit on the bench” rather than just move to its coordinates) and to add or remove objects. As P4-LP pointed out, without supporting object functionality interaction, the objects felt like “dry artifacts” instead of interactive video game objects. Success rates were low, especially for Creation/Deletion, reflecting system limitations rather than lack of interest. However, these attempts highlight aspiration for richer interactivity beyond current system capabilities.
@@ 5.3. Emergent User Strategies and Goals
Exploration was the entry point for nearly all tasks. It included US (understand scene), SS (search in scene), and QI (questions about specific items). In Scene 2, 33 of 48 tasks (68.8%) began with exploration, and in Scene 3, all eight participants opened with such prompts. P5-VI explained: “You have to first ask what is an object, then know what it is, then increase size… and be more specific.” Among the 336 total prompts, QI was most frequent (93), followed by SS (48) and US (36). These patterns highlight the central role of exploration in helping BLV users build an initial mental model of the environment.
@@
After establishing orientation, participants used execution prompts to act on the scene. Codes included EK (external knowledge), EM (explicit modification), PM (proactive modification), and CM (creative modification). EK was rare (7 prompts) but notable, as it showed users using the LLM to probe for accessibility knowledge beyond the scene itself, such as “If I had a camera, would the light sources be bright enough to take a clear picture?” P4-LP highlighted its promise: “you could have the game intuit its own design and say how might this be more accessible in a non-visual way?”
EM and PM were common, aligning with task completion and proactive adjustments. Interestingly, CM appeared in 34 prompts, with participants using the system creatively, such as “Create a door in front of me” (P2-VI) or “add a canopy and luxury padding to the bench” (P7-B). For some, creative use became dominant — P2-VI devoted 10 of her 17 free-exploration prompts to CM. These findings suggest that beyond accessibility, participants viewed the system as a potential authoring tool.
@@ 6. Discussion
Building on these results, we discuss opportunities for safeguarding and verification, the potential role of LLMs as accessibility experts, ways to reduce developer effort through metadata automation, and broader implications for conversational programming. We conclude with limitations and directions for future work.
@@ 6.3. Reducing Developer Burden through Metadata Automation
Our system required manual developer setup to populate the semantic scene graph with object and sound descriptions, limiting scalability. Prior work shows that captioning 3D models and sound files can be partially automated. For instance, Cap3D provides scalable 3D object captioning (Luo et al., 2023), and ExCap3D supports expressive, variable-level descriptions (Yeshwanth et al., 2025). Audio language models (e.g., Audio Flamingo (Kong et al., 2024)) can generate captions for sound sources.

### v2

@@ 4.3.1. Participants
4.3.2. Study Setup
The scenes in the study were deployed in a 3D environment running on a laptop. Prior work shows that this setup could be effectively used to evaluate visual accessibility in 3D and 360 environments (Nair et al., 2021; Froehlich et al., 2025). We excluded more novel formats (e.g., VR) to keep the study centered on 3D spatial aspect and to avoid confounding the results with extra learning demands. Participants used a laptop device running Unity, a full-sized keyboard, and stereo headphones. The researcher also used stereo headphones to monitor system audio output in real time. See fig. 4.
The image shows the study setup. A BLV participant and a researcher are sitting at a table with a laptop. The laptop screen is showing Scene 2 with RAVEN in action. The BLV user is using a keyboard to interact with the scene. Both the participant and the researcher had stereo headphones to access the sounds from the scene.
Figure 4. The study setup for the evaluation of RAVEN with BLV participants.
@@ 4.3.3. Procedure
Studies lasted approximately 1.5 hours and were conducted either in-lab (N=6) or at participants’ homes (N=2). Each session followed the three scene explorations (section 4.2), then a survey (usefulness of categories, confidence, intuitiveness, and SUS (Brooke, 1996)), and finally a semi-structured interview (usability, new category ideas, prompt strategies, accessibility needs, AI trust). Sessions were audio- and video-recorded. Participants received a \$50 gift card and travel reimbursement. The study was IRB-approved.
@@ 4.3.4. Data Analysis
We analyzed three sources of data: survey ratings, prompt logs, and interview transcripts. Together, these provided complementary perspectives on participants’ experiences and interaction strategies. The research team is mixed-ability, with one BLV, one DHH, and three able-bodied researchers. We acknowledge that these backgrounds shape our interpretation.
@@
The survey was designed to address RQ1, evaluating RAVEN’s support for BLV users in 3D scenes. It included an SUS questionnaire (Brooke, 1996) to evaluate system usability, and 5-point Likert scale questions about: usefulness of categories (section 4.1), user confidence in accessibility improvement, and intuitiveness of the system. We computed descriptive statistics (means, standard deviations) to summarize the data.
@@
Prompts and system responses from Scenes 2 and 3 were recorded, yielding 336 valid prompts (181 from Scene 2 and 155 from Scene 3), after excluding the tutorial scene and erroneous inputs. We collected response time from recordings, measuring from the frame of user input submission to the frame when the LLM finished loading and reply appeared. Each prompt was independently coded by multiple researchers along three dimensions, with disagreements resolved through discussion:
@@
Interviews were transcribed and analyzed using applied thematic analysis (Guest et al., 2012). Three researchers collaboratively: (1) familiarized themselves with the transcripts, (2) tagged data with codes, (3) developed an initial codebook, (4) refined codes and themes, (5) finalized the codebook, and (6) synthesized themes into findings.
This process produced a final codebook with 114 codes, organized into 30 third-level themes, 12 second-level themes, and 4 first-level themes. Before integrating qualitative insights into the Findings section, we summarize the structure of our thematic analysis. Our final codebook contained four major themes and twelve subthemes: (1) System Performance and Reliability (perceived accuracy and trust, handling of hallucinations, and verification strategies); (2) Interaction and Feedback Experience (intuitiveness of natural language interaction, appropriation of iterative prompting, and modality preferences across audio and visual cues); (3) Prompt Categories and Use Patterns (task-driven use in Scene 2, open-ended exploration in Scene 3, and differences between effective and ineffective prompting strategies); and (4) Ethical and Broader Impacts (concerns about safety and scene integrity, desires for guardrails and transparency, and reflections on broader accessibility opportunities and risks). These themes structure the qualitative findings reported in section 5 and clarify how we organized perspectives on system performance, interaction experience, prompting behavior, and broader impacts.
@@ 5. User Study Findings
Our analysis highlights both the promise and current limitations of RAVEN. We first report the system’s performance, followed by participants’ overall perceptions of usability, then examine performance across the six guiding and four emerging prompt categories, and finally describe the strategies and goals that shaped how participants engaged with the system.
5.1. System Performance Evaluation
Of the 336 valid prompts from Scenes 2 and 3 in the user study, 253 (75.3%) produced a correct query answer or intended modification (determined through researcher coding of session recordings), 9 (2.7%) were correctly flagged as out-of-scope, and 74 (22.0%) failed. Failures included intent errors (14 prompts, where the LLM misunderstood participant goals) and technical errors (60 prompts, where the LLM hallucinated, misrepresented system capabilities, or failed in code execution).
The average response time was 3.1 seconds (SD=3.7SD{=}3.7) across all prompts. A breakdown of the average response time for different correctness labels is shown in table 3. Acknowledge Out-of-scope prompts had the shortest average response time and the lowest variability. The remaining correctness labels had similar average response times, with Technical Errors requiring slightly longer.
Table 3. Counts and percentages of prompts per correctness label, and their corresponding average response times.
Correctness Label Count Percentage (%) Avg. Response Time (sec) SD of Response Time (sec)
Success 253 75.3 3.1 4.1
Acknowledge Out-of-scope 9 2.7 2.1 0.5
Intent Errors 14 4.2 3.2 1.6
Technical Errors 60 17.9 3.4 2.5
Total Errors (Intent and Technical) 74 22.0 3.2 2.3
All Prompts 336 - 3.1 3.7
@@ 5.2. Overall System Usability
Despite these advantages, several limitations emerged. P4-LP described the gaps in flexible language and cross-modal consistency: extinguishing a torch stopped the light but left the fire sound. P6-VI similarly noted misapplied assumptions, such as turning all objects white after asking to “make the room white”. Participants (N=7) raised trust concerns about misleading responses. As P2-VI asked, “If it gives the wrong information, how is a visually-impaired person going to know?” P4-LP concluded: “right now, [the system] is not so foolproof as to say I completely trust it.” These findings highlight both the promise and current fragility of the approach. Because our participants’ visual abilities spanned a continuum—from no usable vision to varying forms of low vision—these impressions reflect a spectrum of experiences rather than a clean blind/low-vision split. This motivates our choice to report visual-ability differences qualitatively instead of conducting binary group comparisons.
5.3. Performance and Usage Across Prompt Categories
Qualitative coding of 336 prompts revealed ten categories: six guiding categories defined in study design (Object Location, Audio Volume, Color, Object Size, Scene Brightness, Audio Pitch) and four emerging categories that surfaced during analysis (Scene Description, Semantic Description, Functionality, Creation/Deletion). Nine prompts were coded as “Other” and eight combined multiple categories (e.g., “Make the bench normal-sized (Object Size) and put us on the bench” (Object Location)). See table 4 for code definitions and examples.
@@ 5.3. Performance and Usage Across Prompt Categories
The image shows a histogram of the total occurrence of each prompt category, sorted from highest to lowest for guiding categories and for emerging categories. For the guiding categories, from the highest to lowest are Object Location being higher than 70; then Color and Audio Volume at around 40; Object Size between 30-40; Scene Brightness at around 30, Audio Pitch at around 10. For emerging categories, from highest to lowest are Scene Description at around 40, Semantic Description at around 30, Functionality at around 10, and Creation and Deletion at around 5.
(a) Occurrences of each category (blue = guiding, orange = emerging).
The image shows a stacked bar chart of the percentage of correctness codes breakdown by categories. The last column shows the percentage data for all prompts combined.
(b) Correctness breakdown per category.
@@ 5.3.1. Guiding Categories
Lowest rated ($M=2.8,SD=1.7$), Audio Pitch adjustments were less useful overall. Some participants envisioned benefits for distinguishing status (e.g., damage to a game character) or orientation (P7-B: “using pitch for positional information is a really good idea… for example, gets higher when you get closer to it”), but most preferred volume as a more salient cue. Participant comments suggested potential for system-level pitch cues for important events or orientation in future designs, but this category had limited value as a user-driven adjustment in our scenarios.
@@ 5.3.2. Emerging Categories
Although not predefined, these categories were frequently used, revealing additional details preferred by participants. Scene descriptions offered general overviews, while semantic descriptions probed higher-level qualities (e.g., the mood of a cat or the content of a laptop screen). Participants valued detail but noted issues of verbosity: P4-LP suggested limiting output to nearby objects or allowing control over granularity. P7-B emphasized context-aware tailoring: “Maybe all I need to know is… I don’t need to know that it’s a Toshiba.” Out-of-scope acknowledgments were common, as users often sought knowledge beyond system capabilities. Suggested improvements included adding shape (P1-LP), clearance (P7-B), sound characteristics (P2-VI, P8-B), and hazard indicators (P7-B).
@@
Although infrequent, these categories revealed unmet needs. Participants wanted to interact with objects functionally (e.g., “sit on the bench” rather than just move to its coordinates) and to add or remove objects. As P4-LP pointed out, without supporting object functionality interaction, the objects felt like “dry artifacts” instead of interactive video game objects. Success rates were low, especially for Creation/Deletion, reflecting system limitations rather than lack of interest. However, these attempts highlight aspiration for richer interactivity beyond current system capabilities.
5.3.3. Category Usage Comparison Between Scene 2 and Scene 3
Scene 2 asked participants to complete specific tasks aligned with the six guiding categories, whereas Scene 3 encouraged open-ended exploration. Figure 6 shows the percentage of prompts in each category for the two scenes (excluding Compound and Other prompts). Five categories showed notably higher relative usage in Scene 3 than in Scene 2: Object Location, Scene Description, Audio Pitch, Functionality, and Creation/Deletion. These differences suggest two broad exploration patterns in the open-ended scene: spatial exploration (Object Location, Scene Description) and novel feature exploration (Audio Pitch, Functionality, Creation/Deletion). Together, they highlight how participants appropriated RAVEN differently when not guided by predefined tasks.
The image shows the prompt category occurrence frequency by scene. On the x axis are the categories, and the y axis is the percentage of prompt categories. The graph shows percentage values for both Scene 2 and Scene 3.
Figure 6. Occurrence frequency of each category within the categorizable prompts (excluding Compound and Other) in each scene (blue = Scene 2 percentage, orange = Scene 3 percentage).
@@ 5.4. Emergent User Strategies and Goals
Exploration was the entry point for nearly all tasks. It included US (understand scene), SS (search in scene), and QI (questions about specific items). In Scene 2, 33 of 48 tasks (68.8%) began with exploration, and in Scene 3, all eight participants opened with such prompts. P5-VI explained: “You have to first ask what is an object, then know what it is, then increase size… and be more specific.” Among the 336 total prompts, QI was most frequent (93), followed by SS (48) and US (36). These patterns highlight the central role of exploration in helping BLV users build an initial mental model of the environment. The success rate for US prompts (86.1%) was higher than for SS (75.0%) and QI (73.1%), suggesting that strategies starting with broad, whole-scene queries were more reliably supported than strategies that immediately targeted specific objects.
@@
After establishing orientation, participants used execution prompts to act on the scene. Codes included EK (external knowledge), EM (explicit modification), PM (proactive modification), and CM (creative modification). EK was rare (7 prompts) but notable, as it showed users using the LLM to probe for accessibility knowledge beyond the scene itself, such as “If I had a camera, would the light sources be bright enough to take a clear picture?” P4-LP highlighted its promise: “you could have the game intuit its own design and say how might this be more accessible in a non-visual way?” EK prompts also demonstrated a relatively high success rate (85.7%), underscoring that this strategy worked well when participants treated the LLM as an accessibility advisor.
EM and PM were common, aligning with task completion and proactive adjustments. Interestingly, CM appeared in 34 prompts, with participants using the system creatively, such as “Create a door in front of me” (P2-VI) or “add a canopy and luxury padding to the bench” (P7-B). For some, creative use became dominant — P2-VI devoted 10 of her 17 free-exploration prompts to CM. Creative modifications were sometimes successful but also more likely to expose limitations (e.g., unsupported functionality or implausible object behavior), indicating that CM is a powerful but less predictable strategy compared to EM and PM. These findings suggest that beyond accessibility, participants viewed the system as a potential authoring tool.
@@
6. Preliminary Evaluation with Unity Developers
6.1. Method
To understand how RAVEN fits into real development workflows and to evaluate its scalability from a developer perspective, we conducted a preliminary usability study with six Unity developers (three men, three women). Participants were recruited through forum posts and snowball sampling. Their Unity experience ranged from 0.5 to 6 years (M=1.8M{=}1.8, SD=1.9SD{=}1.9). Two participants had prior experience with accessibility toolkits (e.g., Meta XR accessibility tools), and two had used LLMs in game development.
Researchers first demonstrated how to integrate RAVEN into a simple example Unity scene (five objects including two 3D objects, a sound source, a light source, and a text object) following the workflow described in section 3.3.1. Participants then repeated this integration themselves in the example scene. Next, they incorporated RAVEN into one of their own scenes, either self-created or selected from online resources. Written instructions accompanied the implementation steps, and participants were not given time constraints.
We recorded (1) the time required to achieve a working integration in the example scene and (2) the time needed to apply a functioning version of RAVEN to a personal scene. Time was measured from the start of implementation to the first successful runtime test of an LLM prompt. Although we did not explicitly prompt developers to think aloud, several spontaneously verbalized thoughts about the system during the process. Because some personal scenes were large, we also collected estimates of how many objects would require tagging for full coverage.
After implementation, participants completed a questionnaire (5-point Likert-scale items on learnability, usability, perceived accessibility improvement, and intent for future use), followed by a semi-structured interview to capture experiences, preferences, and improvement suggestions. Sessions were conducted in person, lasted approximately one hour, and participants received \$30 compensation. Quantitative ratings were summarized using descriptive statistics. Two coders performed thematic analysis (Guest et al., 2012), producing 12 codes grouped under four themes—initial learning, developer interaction, system behavior, and perceived accessibility impact (see supplementary materials for full code book).
6.2. Findings
Developers spent an average of 265.8 seconds (SD=92.6SD{=}92.6) integrating RAVEN into the example scene and 331.8 seconds (SD=63.2SD{=}63.2) applying a working version to their own scenes. Estimated tagging effort averaged 26.7 objects per scene (SD=16.7SD{=}16.7), though estimates varied with scene complexity and the tagging strategies developers preferred.
Initial Learning and Setup. Developers described the setup process as easy and intuitive. Learnability received a mean rating of 4.7 (SD=0.5SD{=}0.5). All six participants commented on the ease of initial system setup. Three appreciated the drag-and-drop design, while two noted that such simplicity could particularly help novices who may feel overwhelmed by other accessibility tools. A recurring challenge involved unclear variable names (e.g., using isMeta to denote non-physical objects like sunlight); four participants recommended more transparent naming conventions.
Developer Interaction and Tagging Strategies. Usability ratings remained positive when applying RAVEN to developers’ own scenes (M=4.3M{=}4.3, SD=0.5SD{=}0.5). Tagging strategies varied: some developers preferred tagging only interactable objects, while others advocated tagging everything visible for parity with sighted access. D4 noted that tagging hidden or narrative-sensitive elements (e.g., easter eggs) might unintentionally spoil gameplay, whereas D5 suggested including a tagging field to mark the importance of each object. D6 argued that “a sighted person can see all the objects, so to provide an equivalent experience, we should tag all of them.”
System Behavior and Desired Automation. Developers suggested automating several aspects of the workflow, including: generating unique object names (D2), enabling batch tagging (D3), collecting renderers from nested objects (D4), and generating text descriptions from models or images (D6). D2 also suggested an interface for customizing model and prompt settings, though this might increase the learning curve.
Perceived Accessibility Impact and Adoption. Developers rated RAVEN’s potential to improve accessibility highly (M=4.7M{=}4.7, SD=0.5SD{=}0.5) and said they would use it in future projects (MM=4.7, SS=0.5). Five participants felt RAVEN could meaningfully enhance BLV user experience. Interestingly, D3 described the system as a connector between developers and disabled users: developers supply descriptions and structural intent, while the LLM tailors them to user needs.
However, three participants raised concerns about LLM-generated errors, echoing concerns from prior work in runtime generation (Jennings et al., 2024) and feedback from BLV participants. For future improvements, participants suggested features such as event-history tracking (D6), improved input/output methods (N=4), and using RAVEN not only for modifying scenes but also for authoring accessible scenes (N=2).
@@ 7. Discussion
Throughout the study, participants appropriated RAVEN in different ways across Scenes 2 and 3, highlighting both task-driven use and open-ended exploration. These behaviors, together with the error profile we observed, inform how RAVEN should be positioned relative to existing accessibility tools and how future systems might provide safe, scalable support. Building on these results, we discuss opportunities for safeguarding and verification, the potential role of LLMs as accessibility experts, ways to reduce developer effort through metadata automation, and broader implications for conversational programming. We conclude with limitations and directions for future work.
@@ 7.2. LLMs as Accessibility Experts
7.3. Positioning RAVEN Among Accessibility Approaches
RAVEN occupies a distinct space among existing accessibility approaches for virtual 3D environments. Static toolkits such as SeeingVR (Zhao et al., 2019b) offer valuable enhancements for low-vision users but rely on developer-authored overlays and visual filters. These toolkits do not support runtime code generation, natural language prompting, or modification of audio and spatial structures. As a result, they are complementary to rather than directly comparable with RAVEN, which focuses on enabling BLV users to interactively query and modify 3D environments in real time.
An alternative design question is whether large language models could simply “make the entire scene accessible” in a single transformation. Our findings suggest that such one-shot global modifications would not meet BLV users’ needs. Participants expressed diverse and sometimes conflicting accessibility preferences across the visual-ability spectrum—for example, some preferred brighter scenes while others needed dimmer ones, and some relied on color semantics while others valued purely spatial descriptions. Participants also engaged in iterative exploration, targeted adjustments, and verification, using conversational interaction to personalize accessibility in ways that a single global transformation could not capture. Importantly, accessibility needs emerged as highly individualized and context-dependent in our study, suggesting that adaptive, user-guided modification workflows are more appropriate than rigid, global transformations.
By enabling controlled, query-guided modifications grounded in embodied metadata and accessibility rules, RAVEN supports accessible interaction as an iterative, user-directed process. This design preserves user agency, aligns modifications with contextual grounding, and avoids the brittleness of one-shot transformations, positioning RAVEN as a complementary and necessary approach alongside existing static and LLM-based accessibility tools.
@@ 7.4. Reducing Developer Burden through Metadata Automation
Our system required manual developer setup to populate the semantic scene graph with object and sound descriptions (section 6). In our preliminary developer study, participants found the overall workflow learnable and usable, but expressed concern about scaling tagging and description efforts to larger scenes. Several requested automation to generate names, collect renderers from nested objects, and produce first-pass descriptions for both visual and audio elements.
These requested capabilities can be supported in future integrations of RAVEN’s authoring into mainstream developer environments. For example, Cap3D provides scalable 3D object captioning (Luo et al., 2023), and ExCap3D supports expressive, variable-level descriptions (Yeshwanth et al., 2025). Audio language models (e.g., Audio Flamingo (Kong et al., 2024)) can generate captions for sound sources. Accessibility-focused plugins such as UI Accessibility Plugin (Metalpop Games, 2025) could integrate these intelligent capabilities to generate a first-pass layer of accessibility metadata for developers.